MILLIMETERWAVE SPECTRUM OF ETHYLENE

\underline{GEORGES WLODARCZAK}, JEAN DEMAISON, LAURENT MARGULES, BILAL BAKRI, and FRANCOIS HERLEMONT , Laboratoire de Physique des Lasers, Atomes et Molécules, UMR CNRS 8523, Université de Lille 1, F-59655 Villeneuve d’Ascq, France ; ANDRE FAYT , Molecular Spectroscopy Laboratory, Université Catholique de Louvain, Chemin du Cyclotron 2, B-1348 Louvain-La-Neuve, Belgium.

The rotational spectra of deuterated isotopomers of ethylene, H_{2}C=CH_{2}, have already been measured ^{a}. The parent species in its ground vibrational state, as far as it is concerned, has no rotational spectrum because it has no permanent dipole moment. However, the vibrational states v_{7} = 1 and v_{8} = 1 are close in energy (at 948.77 and 939.86 cm^{−1}, respectively) and some allowed transitions between these two states lie in the millimeterwave range. The ν_{7} state is known with an accuracy better than 10^{−6} cm^{−1} thanks to CO_{2} laser sideband spectra ^{b}. The ν_{8} state has been estimated with an accuracy of the order of 10^{−3} cm^{−1} by combining the ν_{7} + ν_{8} band with the hot band ν_{7} !+ ν_{8} − ν_{8}, both analysed in F.T. spectra (unpublished results). The spectrum has been measured from 180 to 500 GHz and more than 30 transitions have been assigned without ambiguity to ethylene, with a systematic deviation of about -27 MHz from the predictions. A new analysis of the ν_{8} state has been performed, including the a-type Coriolis interaction with the ν_{6} state, and a good agreement has been obtained.

aE. Hirota, Y. Endo, S. Saito, K. Yoshida, I. Yamaguchi, and K. Machida, J. Mol. Spectrosc. 89 (1981) 223.

bE. Rusinek, H. Fichoux, M. Khelkhal, F. Herlemont, J. Legrand, and A. Fayt, J. Mol. Spectrosc. 189, 64-73 (1998).